OpenStudy (anonymous):

What value is a discontinuity of x squared plus 8 x plus 4, all over x squared minus x minus 6? @antonio_xx2 @e.mccormick @quickstudent @FibonacciChick666 @myininaya @help123456 @helpmeplzzzzz! @dan815 @phi @Pi123

OpenStudy (abb0t):

@dan815

OpenStudy (anonymous):

\[\frac{ x^{2} + 8x + 4 }{ x^{2} - x - 6 }\] This will help @dan815 @abb0t

OpenStudy (fibonaccichick666):

so what value(s) of x make the bottom go haywire?

OpenStudy (abb0t):

factor.

OpenStudy (fibonaccichick666):

but yes, you factor to get to that point

OpenStudy (anonymous):

I factor the equation on top it is hard please help

OpenStudy (fibonaccichick666):

just factor the bottom to strat

OpenStudy (anonymous):

I did and I got (x - 3) (x + 2)

OpenStudy (fibonaccichick666):

Good!

OpenStudy (fibonaccichick666):

now, when does that make the bottom a problem? for which x values?

OpenStudy (anonymous):

But what is the right factors for the top

OpenStudy (pi123):

im sorry but im bad at those kind of problems, so i dont think ill be able to tell you. goodluck though!

OpenStudy (phi):

the top does not factor nicely. so you don't have to factor at all (it matters only if its factors match one of those in the bottom)

OpenStudy (fibonaccichick666):

for the top, you want something that multiplies to 4 but adds to 8, which isn't possible so you have to complete the square, but you don't need the top here

OpenStudy (anonymous):

So what is the answer to my question @FibonacciChick666 ?

OpenStudy (phi):

any x value that causes a divide by 0

OpenStudy (anonymous):

It is asking What value is a discontinuity vaule

OpenStudy (fibonaccichick666):

things that make the bottom go haywire ie =0

OpenStudy (fibonaccichick666):

so, what makes the bottom =0?

OpenStudy (anonymous):

So the answer would be x = -2 or 2

OpenStudy (fibonaccichick666):

put that in and see if it works, does it?

OpenStudy (phi):

how did you get +2 ?

OpenStudy (phi):

here is a graph, to show what's going on

OpenStudy (anonymous):

What value is a discontinuity of x squared plus 8 x plus 4, all over x squared minus x minus 6? x = -1 x = -4 x = -3 x = -2 I don't know This is the whole question @FibonacciChick666 @phi

OpenStudy (fibonaccichick666):

well, check it for us then. Input the x value and see if you get zero or not

OpenStudy (phi):

you found ***I did and I got (x - 3) (x + 2)*** what would make (x - 3) (x + 2) =0 ?

OpenStudy (phi):

do you see if x is -2, then (x - 3) (x + 2) would be (-2-3) (-2 + 2) = -5*0 = 0 or if x = 3 you would get (3-3)(3+2) = 0 * 5 = 0 so those are your discontinuities

OpenStudy (anonymous):

So the answer is -2 right?

OpenStudy (phi):

to solve (x - 3) (x + 2) =0 you make it two separate problems x-3= 0 and x+2 =0 we get x= 3 and x = -2 (if you know how to solve these equations)

OpenStudy (phi):

There are two answers. But you only have to pick the one they listed.

OpenStudy (anonymous):

-2 is the right answer @phi

OpenStudy (fibonaccichick666):

is -2 one of the answers?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thanks for the help but, I can only give one person a medal and I will Fan the other